# Higher-Order Regularization Learning on Hypergraphs

### v1

@@ 1 Introduction
Graphs play a foundational role in machine learning, enabling effective modeling of relational data across a range of tasks—from semi-supervised learning and clustering to recommendation systems and manifold learning, e.g. [96, 94, 95, 55, 11, 26, 57, 88, 48]. However, many real-world phenomena involve more complex interactions among sets of nodes, that are not fully captured by pairwise edges. Hypergraphs extend graphs by allowing hyperedges to connect arbitrary subsets of nodes, and hypergraph-based methods are used broadly in various areas of science such as in [30, 90, 67, 44, 71, 65, 17, 29, 92, 60, 93, 54, 24].
@@
A central research question concerns the comparison between hypergraph and graph-based learning methods. Many such comparisons are grounded in discrete arguments (e.g.,[1, 46, 18, 47, 59]). More recently, asymptotic consistency frameworks—a popular technique for analyzing graph-based methods by relating discrete energies to continuum variational limits (e.g.,[4, 19, 39, 73, 78, 9, 75, 22, 87])—have been extended to the hypergraph regularization setting [71, 85]. This continuum perspective allows for a principled assessment of the role of hypergraph structures, supports a classification of hypergraph learning algorithms [85, Figure 2], and enables a clearer understanding of the regularization behavior underlying complex discrete formulations.
@@
In the analysis of graph- and hypergraph-based regularization, it is useful to distinguish between two complementary components of a regularizer: (1) the support of interactions, i.e., which nodes influence one another (determined by the graph or hypergraph topology), and (2) the interaction mechanism, i.e., how these influences are aggregated or penalized (e.g., via first-order differences, higher-order derivatives, or more general nonlinear terms). Classical hypergraph learning methods typically enrich the interaction support—by allowing edges to connect sets of nodes rather than pairs—but still rely on first-order, pairwise-like regularization mechanisms [93, 85].
@@
In this context, Higher-Order Hypergraph Learning (HOHL) was introduced as a method that more effectively leverages hypergraph structure—not only by modifying which interactions are considered, but also by altering the nature of such interactions. Specifically, HOHL decomposes the hypergraph into a sequence of subgraphs that capture interactions at multiple scales. On each subgraph, a distinct regularization strength is applied, allowing the model to enforce higher-order smoothness in a structured and scale-aware manner. In doing so, HOHL exploits the full expressive potential of the hypergraph more fully and effectively. From an analytical perspective, HOHL is shown to converge to a higher-order Sobolev semi-norm, making it genuinely distinct from other hypergraph methods [93, 71] that asymptotically recover the standard $\mathrm{W}^{1,p}$ regularization.
@@ 1.2 Related works
A growing body of work has focused on the asymptotic consistency and continuum analysis of graph-based regularization in the large-sample regime. These efforts include convergence results for total variation on graphs [35], graph cuts and Cheeger-type problems [38, 37, 33, 62], the Mumford–Shah functional [15], and empirical risk minimization [32]. In the semi-supervised setting, particular attention has been given to $p$-Laplacian learning [75], fractional Laplacian methods [87], Lipschitz learning [51, 10, 66, 8], game-theoretic formulations [9], Poisson learning [11, 7], reweighted Laplacians [72], and truncated energy models [2, 3]. These developments reflect a general trend toward understanding the behavior of discrete algorithms through the lens of continuum variational principles. Recently, such analyses have been extended to the hypergraph setting [71, 85]
@@
  • Pointwise convergence [4, 19, 39, 43, 42, 73, 78] examines whether $\mathcal{E}_{n}(v|_{\Omega_{n}})\to\mathcal{E}_{\infty}(v)$ as $n\to\infty$, for sufficiently smooth functions $v:\Omega\to\mathbb{R}$. A related approach considers the pointwise convergence of the associated Euler–Lagrange operators [86].
@@
  • Spectral convergence [4, 31, 12, 83, 63, 74, 84] analyzes the convergence of the eigenvalues and eigenfunctions of the discrete operator associated with (through Euler-Lagrange equations) $\mathcal{E}_{n}$ to those of the limiting operator appearing in $\mathcal{E}_{\infty}$.
@@
  • Variational convergence [9, 20, 22, 35, 38, 37, 36, 75, 77, 80] concerns the convergence of minimizers of $\mathcal{E}_{n}$ to those of $\mathcal{E}_{\infty}$, typically formalized through $\Gamma$-convergence [6]. Among the three notions, it is often the most relevant in semi-supervised learning, where the final label assignments are derived from minimizers of the objective functional.
@@
In this work, we focus on the latter two modes of convergence. In particular, to establish the variational convergence of our truncated energies, we analyze the spectral properties of the HOHL Laplacian [85]. Our results in the fully supervised setting are also of variational type, providing convergence guarantees for minimizers of the discrete energies.
@@
While much of the literature has focused on consistency, in the graph-based setting, recent works established convergence rates in terms of various parameters such as the number of points $n$, the labeling rate, the graph connectivity parameter $\varepsilon$ and the smoothness of the target function [86, 14, 23]. These rates offer important theoretical guarantees for practical applications, where the dataset is finite and the discrete approximation error must be controlled. In this work, we extend such results to the HOHL framework, similarly to [34], showing explicit convergence rates between the discrete minimizers and the continuum ground truth under suitable regularity assumptions.
@@
Beyond rates, computational efficiency is a key concern for applications. In practice, Laplace learning and related graph-based methods often rely on a spectrally truncated energy formulation. While these truncations are computationally efficient and widely adopted in large-scale settings, their theoretical justification has largely remained heuristic [2, 5, 58]. In this work, we contribute to closing this gap by showing that even when the HOHL energy is truncated, it remains variationally consistent with the full model and converges to the same continuum limit.
@@
Our final computational result establishes a connection between the HOHL Laplacian, and a broad body of work on graph reweighting [72, 13] (in classical models) and graph rewiring[79, 40, 61, 49, 25, 50] (in graph neural networks), both of which aim to improve learning performance by structurally modifying the graph. These modifications are often employed to address limitations such as oversmoothing or oversquashing [40]. In the spirit of [1], which advocates for representing hypergraph structure within enriched graph formulations, we show that HOHL can be interpreted as Laplace learning on a modified graph constructed directly from the original hypergraph structure.
@@ 2 Background
This section presents the mathematical tools used throughout the paper. We begin by recalling the $\mathrm{TL}^{p}$ space, which provides a natural topology for comparing functions defined on discrete empirical measures to functions on the continuum. We then review key concepts from $\Gamma$-convergence theory, which we rely on to study the asymptotic behavior of our discrete variational problems. References for the material presented here include [35, 75, 87, 6].
@@ 2.1 The $\mathrm{TL}^{p}$ Topology
This framework allows us to treat discrete functions—defined on sampled data—as elements of a well-defined metric space and to compare them to their continuum counterparts in a stable way. The topology is closely related to the $p$-Wasserstein [81, 68] on the graph of the function.
A useful characterization of convergence in $\mathrm{TL}^{p}$ is the following [35, Proposition 3.12].
@@
To apply this result, we rely on the following result (see [31, Theorem 2] or [87, Theorem 2.3]) establishing that such transport maps exist for empirical measures constructed from i.i.d. samples.
@@ 2.2 $\Gamma$-Convergence of Functionals
Suppose that $F_{n}:Z\mapsto[0,\infty]$ $\Gamma$-converge to $F:Z\mapsto[0,\infty]$. Furthermore, assume that $G:Z\mapsto[0,\infty]$ is continuous. Then, $F_{n}+G$ $\Gamma$-converge to $F+G$.
@@ 3 Main results
In this section, we present our main results as well the relevant notation and assumptions used for our proofs.
@@ 3.1 Hypergraphs
Figure 1: From graphs to hypergraphs (from [85]). Left: In the graph, the vertices $v_{1}$, $v_{2}$, and $v_{3}$ are all connected pairwise. Right: A single hyperedge is added connecting all three vertices, transitioning from a graph to a hypergraph representation.
@@
Figure 2: Skeleton graphs with $q=2$ (from [85]).
@@ 3.2 HOHL
On graphs, a widely used regularizer is constructed using the graph Laplacian [96, 82]. For a function $u:V\to\mathbb{R}$ (which we also identify with a vector in $\mathbb{R}^{|V|}$), its first-order smoothness is quantified by
@@
Minimizing this expression encourages $u$ to take similar values on adjacent vertices. On certain graphs, this functional can be interpreted as a discrete analogue of the Sobolev $\mathrm{W}^{1,2}$ semi-norm, which formalizes the idea of penalizing the first derivative of a function defined on the graph [75]. More generally, the regularizer $v^{\top}L^{s}v$, with $s\in\mathbb{R}$, corresponds to a discrete Sobolev $\mathrm{W}^{s,2}$ semi-norm and penalizes variations of $v$ up to order $s$ [22, 87].
The HOHL energy, introduced in [85], extends graph Laplacian regularization to the hypergraph setting. It is defined as
@@
In the graph setting, this idea is typically implemented via $k$-nearest neighbor ($k$-NN) graphs [82] or random geometric graphs [64], both of which rely on locality: edges are formed either by linking the $k$ nearest neighbors or by connecting points within an $\varepsilon$-radius neighborhood. Analogous locality-based constructions for hypergraphs have been proposed, e.g., in [71]; see also [30] for a broader discussion. A notable instance is also the random geometric hypergraph model introduced in [85].
As established in [85], the hierarchical, scale-aware regularization principle underlying HOHL admits an effective surrogate in geometric settings via a multiscale graph construction, as proposed in [57]. In what follows, we introduce this alternative formulation.
Let $\Omega_{n}=\{x_{i}\}_{i=1}^{n}\subset\Omega\subset\mathbb{R}^{d}$ be a set of $n$ feature vectors, where we assume that $x_{i}\overset{\text{i.i.d.}}{\sim}\mu\in\mathcal{P}(\Omega)$. We adopt the same probabilistic framework as in [87]. Specifically, we consider a probability space $(\Omega,\mathbb{P})$ whose elements are infinite sequences $\{x_{i}\}_{i=1}^{\infty}$. Our results are stated in terms of the measure $\mathbb{P}$, establishing that the desired properties hold on a high-probability subset $\mathcal{X}\subset\Omega$ consisting of such sequences. For a set $E$, we denote its complement by $E^{c}$.
@@
We note that this is the rescaled version of $L$, i.e. $\Delta_{n,\varepsilon}=\frac{2}{\sigma_{\eta}n\varepsilon^{d+2}}L$. This Laplacian can be interpreted either as a matrix $\Delta_{n,\varepsilon}\in\mathbb{R}^{n\times n}$ or as an operator $\Delta_{n,\varepsilon}:\mathrm{L}^{2}(\mu_{n})\to\mathrm{L}^{2}(\mu_{n})$, where $\mu_{n}=\frac{1}{n}\sum_{i=1}^{n}\delta_{x_{i}}$ is the empirical measure.
@@
where $\varepsilon^{(1)}>\cdots>\varepsilon^{(q)}$, and $p_{k}>0$ controls the regularity imposed at each scale. We will allow the length-scales to vary with the number of data point, i.e. $\varepsilon^{(k)}=\varepsilon_{n}^{(k)}$, and in this case, we write $E_{n}:=\{\varepsilon_{n}^{(k)}\}_{k=1}^{q}$. The well and ill-posedness of (3) in semi-supervised learning is precisely characterized in [85, Theorem 3.5] as a function of $\varepsilon_{n}^{(q)}$.
@@
$\Delta_{\rho}u(x)=-\frac{1}{\rho(x)}\mathrm{div}(\rho^{2}\nabla u)(x),\,x\in\Omega\quad\quad\quad\frac{\partial u}{\partial n}=0,\,x\in\partial\Omega$
and let $\{(\beta_{i},\psi_{i})\}_{i=1}^{\infty}$ be its associated eigenpairs where $\beta_{1}=0<\beta_{2}\leq\beta_{3}\leq\ldots$. Here $\rho$ denotes the density of $\mu$ with respect to Lebesgue measure. We note that $\{\psi_{i}\}_{i=1}^{\infty}$ form a basis of $\mathrm{L}^{2}(\mu)$ and also define
@@
The space $\mathcal{H}^{s}(\Omega)$ is closely related to the Sobolev space $\mathrm{W}^{s,2}(\Omega)$ [22, Lemma 17].
Fully supervised problem with HOHL regularization
@@
We are mainly interested in the case of noisy labels, i.e. when for some $g\in\mathrm{C}^{0}(\Omega)$, we have labels $\mathbf{y}_{n}=\{y_{i}\}_{i=1}^{n}$ where $y_{i}=g(x_{i})+\xi_{i}$ and $\xi_{i}\in\mathbb{R}$ are independent and identically distributed sub-Gaussian centered noise.
Truncated HOHL energies

### v2

@@ 1 Introduction
Graphs play a foundational role in machine learning, enabling effective modeling of relational data across a range of tasks—from semi-supervised learning and clustering to recommendation systems and manifold learning, e.g. [LapRef, zhou2004lazy, zhou2011semi, Mai, poissonLearning, flores2019algorithms, Merkurjev, xia2021graph, ju2024survey]. However, many real-world phenomena involve more complex interactions among sets of nodes, that are not fully captured by pairwise edges. Hypergraphs extend graphs by allowing hyperedges to connect arbitrary subsets of nodes, and hypergraph-based methods are used broadly in various areas of science such as in [hgLearningPractice, clique, hgPLaplacianGeometric, TVHg, shi2025hypergraphplaplacianequationsdata, Pirvu_2023_ICCV, ChenHg, hgWeighting, dynamicHG, neuhauser, scholkopfHyper2006, pmlr-v80-li18e, fazeny].
@@
A central research question concerns the comparison between hypergraph and graph-based learning methods. Many such comparisons are grounded in discrete arguments (e.g.,[hypergraphGraph, jostMulas, chitra, jost, mulas]). More recently, asymptotic consistency frameworks—a popular technique for analyzing graph-based methods by relating discrete energies to continuum variational limits (e.g.,[NIPS2006_5848ad95, COIFMAN20065, Gine, Singer, 10.5555/3104322.3104459, calderGameTheoretic, Slepcev, Stuart, weihs2023consistency])—have been extended to the hypergraph regularization setting [shi2025hypergraphplaplacianequationsdata, weihs2025Hypergraphs]. This continuum perspective allows for a principled assessment of the role of hypergraph structures, supports a classification of hypergraph learning algorithms [weihs2025Hypergraphs, Figure 2], and enables a clearer understanding of the regularization behavior underlying complex discrete formulations.
@@
In the analysis of graph- and hypergraph-based regularization, it is useful to distinguish between two complementary components of a regularizer: (1) the support of interactions, i.e., which nodes influence one another (determined by the graph or hypergraph topology), and (2) the interaction mechanism, i.e., how these influences are aggregated or penalized (e.g., via first-order differences, higher-order derivatives, or more general nonlinear terms). Classical hypergraph learning methods typically enrich the interaction support—by allowing edges to connect sets of nodes rather than pairs—but still rely on first-order, pairwise-like regularization mechanisms [scholkopfHyper2006, weihs2025Hypergraphs].
@@
In this context, Higher-Order Hypergraph Learning (HOHL) was introduced as a method that more effectively leverages hypergraph structure—not only by modifying which interactions are considered, but also by altering the nature of such interactions. Specifically, HOHL decomposes the hypergraph into a sequence of subgraphs that capture interactions at multiple scales. On each subgraph, a distinct regularization strength is applied, allowing the model to enforce higher-order smoothness in a structured and scale-aware manner. In doing so, HOHL exploits the full expressive potential of the hypergraph more fully and effectively. From an analytical perspective, HOHL is shown to converge to a higher-order Sobolev semi-norm, making it genuinely distinct from other hypergraph methods [scholkopfHyper2006, shi2025hypergraphplaplacianequationsdata] that asymptotically recover the standard $\mathrm{W}^{1,p}$ regularization.
@@ 1.2 Related works
A growing body of work has focused on the asymptotic consistency and continuum analysis of graph-based regularization in the large-sample regime. These efforts include convergence results for total variation on graphs [Trillos3], graph cuts and Cheeger-type problems [JMLR:v17:14-490, trillos2017estimating, thorpeCheeger, doi:10.1137/16M1098309], the Mumford–Shah functional [Caroccia_2020], and empirical risk minimization [garcia_trillos_murray_2017]. In the semi-supervised setting, particular attention has been given to $p$-Laplace learning [Slepcev], fractional Laplacian methods [weihs2023consistency], Lipschitz learning [pmlr-v40-Kyng15, doi:10.1137/18M1199241, Bungert, doi.org/10.48550/arxiv.2111.12370], game-theoretic formulations [calderGameTheoretic], Poisson learning [poissonLearning, bungert2024convergenceratespoissonlearning], reweighted Laplacians [shi2017weighted], and truncated energy models [Belkin2002UsingMS, sslManifolds]. These developments reflect a general trend toward understanding the behavior of discrete algorithms through the lens of continuum variational principles. Recently, such analyses have been extended to the hypergraph setting [shi2025hypergraphplaplacianequationsdata, weihs2025Hypergraphs]
@@
  • Pointwise convergence [NIPS2006_5848ad95, COIFMAN20065, Gine, 10.1007/11503415_32, 10.1007/11776420_7, Singer, 10.5555/3104322.3104459] examines whether $\mathcal{E}_{n}(v|_{\Omega_{n}})\to\mathcal{E}_{\infty}(v)$ as $n\to\infty$, for sufficiently smooth functions $v:\Omega\to\mathbb{R}$. A related approach considers the pointwise convergence of the associated Euler–Lagrange operators [weihs2023discreteToContinuum].
@@
  • Spectral convergence [NIPS2006_5848ad95, Trillos, CALDER2022123, 10.1214/009053607000000640, JMLR:v12:pelletier11a, 10.1093/imaiai/iaw016, https://doi.org/10.48550/arxiv.1510.08110] analyzes the convergence of the eigenvalues and eigenfunctions of the discrete operator associated with (through Euler-Lagrange equations) $\mathcal{E}_{n}$ to those of the limiting operator appearing in $\mathcal{E}_{\infty}$.
@@
  • Variational convergence [calderGameTheoretic, cristoferi_thorpe_2020, Stuart, Trillos3, JMLR:v17:14-490, trillos2017estimating, GARCIATRILLOS2018239, Slepcev, thorpe_theil_2019, Gennip] concerns the convergence of minimizers of $\mathcal{E}_{n}$ to those of $\mathcal{E}_{\infty}$, typically formalized through $\Gamma$-convergence [gammaConvergence]. Among the three notions, it is often the most relevant in semi-supervised learning, where the final label assignments are derived from minimizers of the objective functional.
@@
In this work, we focus on the latter two modes of convergence. In particular, to establish the variational convergence of our truncated energies, we analyze the spectral properties of the HOHL Laplacian [weihs2025Hypergraphs]. Our results in the fully supervised setting are also of variational type, providing convergence guarantees for minimizers of the discrete energies.
@@
While much of the literature has focused on consistency, in the graph-based setting, recent works established convergence rates in terms of various parameters such as the number of points $n$, the labeling rate, the graph connectivity parameter $\varepsilon$ and the smoothness of the target function [weihs2023discreteToContinuum, calder2020rates, ElBouchairi]. These rates offer important theoretical guarantees for practical applications, where the dataset is finite and the discrete approximation error must be controlled. In this work, we extend such results to the HOHL framework, similarly to [trillos2022rates], showing explicit convergence rates between the discrete minimizers and the continuum ground truth under suitable regularity assumptions.
@@
Beyond rates, computational efficiency is a key concern for applications. In practice, Laplace learning and related graph-based methods often rely on a spectrally truncated energy formulation. While these truncations are computationally efficient and widely adopted in large-scale settings, their theoretical justification has largely remained heuristic [Belkin2002UsingMS, bertozziStuart, Miller]. In this work, we contribute to closing this gap by showing that even when the HOHL energy is truncated, it remains variationally consistent with the full model and converges to the same continuum limit.
@@
Our final computational result establishes a connection between the HOHL Laplacian, and a broad body of work on graph reweighting [shi2017weighted, properly] (in classical models) and graph rewiring[ricciBronstein, overSmoothing, ricciOsher, rewiringMontufar, rewiringWeber, DGM] (in graph neural networks), both of which aim to improve learning performance by structurally modifying the graph. These modifications are often employed to address limitations such as oversmoothing or oversquashing [overSmoothing]. In the spirit of [hypergraphGraph], which advocates for representing hypergraph structure within enriched graph formulations, we show that HOHL can be interpreted as Laplace learning on a modified graph constructed directly from the original hypergraph structure.
@@ 2 Background
This section presents the mathematical tools used throughout the paper. We begin by recalling the $\mathrm{TL}^{p}$ space, which provides a natural topology for comparing functions defined on discrete empirical measures to functions on the continuum. We then review key concepts from $\Gamma$-convergence theory, which we rely on to study the asymptotic behavior of our discrete variational problems. References for the material presented here include [Trillos3, Slepcev, weihs2023consistency, gammaConvergence].
@@ 2.1 The $\mathrm{TL}^{p}$ Topology
This framework allows us to treat discrete functions—defined on sampled data—as elements of a well-defined metric space and to compare them to their continuum counterparts in a stable way. The topology is closely related to the $p$-Wasserstein distance [villani2009, Santambrogio] on the graph of the function.
A useful characterization of convergence in $\mathrm{TL}^{p}$ is the following [Trillos3, Proposition 3.12].
@@
To apply this result, we rely on the following result (see [Trillos, Theorem 2] based on [garcia_trillos_slepcev_2015]) establishing that such transport maps exist for empirical measures constructed from i.i.d. samples.
@@ 2.2 $\Gamma$-Convergence of Functionals
Suppose that $F_{n}:Z\mapsto[0,\infty]$ $\Gamma$-converge to $F:Z\mapsto[0,\infty]$. Furthermore, assume that $G:Z\mapsto[0,\infty]$ is continuous. Then, $F_{n}+G$ $\Gamma$-converges to $F+G$.
@@ 3 Main results
In this section, we present our main results as well as the relevant notation and assumptions used for our proofs.
@@ 3.1 Hypergraphs
Figure 1: From graphs to hypergraphs (from [weihs2025Hypergraphs]). Left: In the graph, the vertices $v_{1}$, $v_{2}$, and $v_{3}$ are all connected pairwise. Right: A single hyperedge is added connecting all three vertices, transitioning from a graph to a hypergraph representation.
@@
Figure 2: Skeleton graphs with $q=2$ (from [weihs2025Hypergraphs]).
@@ 3.2 HOHL
On graphs, a widely used regularizer is constructed using the graph Laplacian [LapRef, TutSpec]. For a function $u:V\to\mathbb{R}$ (which we also identify with a vector in $\mathbb{R}^{|V|}$), its first-order smoothness is quantified by
@@
Minimizing this expression encourages $u$ to take similar values on adjacent vertices. On certain graphs, this functional can be interpreted as a discrete analogue of the Sobolev $\mathrm{W}^{1,2}$ semi-norm, which formalizes the idea of penalizing the first derivative of a function defined on the graph [Slepcev]. More generally, the regularizer $v^{\top}L^{s}v$, with $s\in\mathbb{R}$, corresponds to a discrete Sobolev $\mathrm{W}^{s,2}$ semi-norm and penalizes variations of $v$ up to order $s$ [Stuart, weihs2023consistency].
The HOHL energy, introduced in [weihs2025Hypergraphs], extends graph Laplacian regularization to the hypergraph setting. It is defined as
@@
In the graph setting, this idea is typically implemented via $k$-nearest neighbor ($k$-NN) graphs [TutSpec] or random geometric graphs [DBLP:books/ox/P2003], both of which rely on locality: edges are formed either by linking the $k$ nearest neighbors or by connecting points within an $\varepsilon$-radius neighborhood. Analogous locality-based constructions for hypergraphs have been proposed, e.g., in [shi2025hypergraphplaplacianequationsdata]; see also [hgLearningPractice] for a broader discussion. A notable instance is also the random geometric hypergraph model introduced in [weihs2025Hypergraphs].
As established in [weihs2025Hypergraphs], the hierarchical, scale-aware regularization principle underlying HOHL admits an effective surrogate in geometric settings via a multiscale graph construction, as proposed in [Merkurjev]. In what follows, we introduce this alternative formulation.
Let $\Omega_{n}=\{x_{i}\}_{i=1}^{n}\subset\Omega\subset\mathbb{R}^{d}$ be a set of $n$ feature vectors, where we assume that $x_{i}\overset{\text{i.i.d.}}{\sim}\mu\in\mathcal{P}(\Omega)$. We adopt the same probabilistic framework as in [weihs2023consistency]. Specifically, we consider a probability space $(\Omega,\mathbb{P})$ whose elements are infinite sequences $\{x_{i}\}_{i=1}^{\infty}$. Our results are stated in terms of the measure $\mathbb{P}$, establishing that the desired properties hold on a high-probability subset $\mathcal{X}\subset\Omega$ consisting of such sequences. For a set $E$, we denote its complement by $E^{c}$.
@@
We note that this is the rescaled version of $L$, i.e. $\Delta_{n,\varepsilon}=\frac{2}{\sigma_{\eta}n\varepsilon^{d+2}}L$. With an abuse of notation, this Laplacian can be interpreted either as a matrix $\Delta_{n,\varepsilon}\in\mathbb{R}^{n\times n}$ or as an operator $\Delta_{n,\varepsilon}:\mathrm{L}^{2}(\mu_{n})\to\mathrm{L}^{2}(\mu_{n})$, where $\mu_{n}=\frac{1}{n}\sum_{i=1}^{n}\delta_{x_{i}}$ is the empirical measure.
@@
where $\varepsilon^{(1)}>\cdots>\varepsilon^{(q)}$, and $p_{k}>0$ controls the regularity imposed at each scale. We will allow the length-scales to vary with the number of data point, i.e. $\varepsilon^{(k)}=\varepsilon_{n}^{(k)}$, and in this case, we write $E_{n}:=\{\varepsilon_{n}^{(k)}\}_{k=1}^{q}$. The well and ill-posedness of (3) in semi-supervised learning is precisely characterized in [weihs2025Hypergraphs, Theorem 3.5] as a function of $\varepsilon_{n}^{(q)}$.
@@
$\Delta_{\rho}u(x)=-\frac{1}{\rho(x)}\mathrm{div}(\rho^{2}\nabla u)(x),\,x\in\Omega$
and let $\{(\beta_{i},\psi_{i})\}_{i=1}^{\infty}$ be its associated eigenpairs where $\beta_{1}=0<\beta_{2}\leq\beta_{3}\leq\ldots$. Here $\rho$ denotes the density of $\mu$ with respect to Lebesgue measure. We note that $\{\psi_{i}\}_{i=1}^{\infty}$ form a basis of $\mathrm{L}^{2}(\mu)$ and we also define
@@
The space $\mathcal{H}^{s}(\Omega)$ is closely related to the Sobolev space $\mathrm{W}^{s,2}(\Omega)$ [Stuart, Lemma 17].
Fully supervised problem with HOHL regularization.
@@
We are mainly interested in the case of noisy labels, i.e. when for some $g\in\mathrm{C}^{0}(\Omega)$, we have labels $\mathbf{y}_{n}=\{y_{i}\}_{i=1}^{n}$ where $y_{i}=g(x_{i})+\xi_{i}$ and $\xi_{i}\in\mathbb{R}$ are independent and identically distributed sub-Gaussian centered noise. Adopting the terminology of [Cucker], for $u_{n,\tau}^{(\mathbf{y}_{n})}$ and $u_{\tau}$ the minimizers of $\mathcal{R}_{n,\tau}^{(\mathbf{y}_{n})}$ and $\mathcal{R}_{\infty,\tau}^{(g)}$ respectively, we regard $u_{n,\tau}^{(\mathbf{y}_{n})}$ as an estimator of $g$. We further decompose its error into a variance and bias component, defined as
$\|u_{n,\tau}^{(\mathbf{y}_{n})}-u_{\tau}|_{\Omega_{n}}\|_{\mathrm{L}^{2}(\mu_{n})}\qquad\qquad\text{and}\qquad\qquad\|u_{\tau}-g\|_{\mathrm{L}^{2}(\mu)}$
respectively. Intuitively, the variance term quantifies the fluctuation caused by finite sampling, while the bias term measures the approximation error introduced by the regularization.
Truncated HOHL energies.